# 2cm2cm1.5cm1.5cm

### v1

@@ 5.2. Inradius in ${F}^{\bot}$
In the centrally symmetric case, we know that the origin is the center of an inball. Hence, the inradius cannot exceed the distance from the origin to a hyperplane $\left\{x\in{F}^{\bot}:\;\left\langle P_{{F}^{\bot}}u_{i},x\right\rangle\leq 1\right\}$ in ${F}^{\bot}$ for some non-zero $P_{{F}^{\bot}}u_{i}.$ This distance is exactly $\frac{1}{\left|P_{{F}^{\bot}}u_{i}\right|}.$ Identities (1.1) and (1.4) imply that $\left|P_{{F}^{\bot}}u_{j}\right|^{2}\geq\frac{d-s}{d}$ for some $j\in[m].$ The bound follows. ∎
@@ 5.3. Volume ratio in $F$
$\int_{\mathbb{R}^{n}}\prod_{i\in[m]}f_{i}(\left\langle x,v_{i}\right\rangle)^{\beta_{i}}\,\mathrm{d}x\leq\prod_{i\in[m]}\left(\int_{\mathbb{R}}f_{i}\right)^{\beta_{i}},$
@@
Let $I\subset[m]$ be the set of those indices for which $P_{F}u_{i}\neq 0$. Set $v_{i}=\frac{P_{F}u_{i}}{\left|P_{F}u_{i}\right|}$ and $\beta_{i}={\upalpha_{i}}{\left|P_{F}u_{i}\right|^{2}}$ for all $i\in I.$ Then
@@
We leave it to the reader to check that the integral on the left-hand side of the Brascamp–Lieb inequality (Lemma 5.6) is equal to the volume of $K_{\mathrm{J_{sym}}}\cap F$, as we identify $F$ with $\mathbb{R}^{n}$ (where $n=s$), and that the product of the integrals on the right-hand side of the Brascamp–Lieb inequality is equal to $2^{s}\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{-\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}.$ The latter quantity is at most $2^{s}\left(\frac{d}{s}\right)^{\frac{s}{2}}$ by routine analysis, completing the proof of Lemma 5.7. ∎
@@
Similarly to Ball’s argument [Bal91], we lift the picture to $\widetilde{F}=F\times\mathbb{R}$ and will write $(a,t)$ for a vector in $\widetilde{F}$ with $a\in F$ and $t\in\mathbb{R}.$
Let $I\subset[m]$ denote the subset of indices $i$ for which $P_{F}u_{i}\neq 0$. Denote $d^{\prime}=\sum\limits_{i\in[m]}\upalpha_{i};$ and set
@@
We will denote the function $t\mapsto e^{-\frac{t}{\left|\widetilde{q}_{i}\right|}}\mathbf{1}_{[0,+\infty)}$ by $f_{i}$ for each $i\in I,$ where $\mathbf{1}_{S}$ stands for the indicator function of the set $S.$ Then, the product of the integrals on the right-hand side of the Brascamp–Lieb inequality (Lemma 5.6) is equal to $\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}.$ The bound
$\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}\leq\frac{(s+1)^{\frac{s+1}{2(d^{\prime}+1)}}}{(d^{\prime}+1)^{\frac{d^{\prime}+1}{2(d^{\prime}+1)}}}$
follows from the standard majorization argument for the convex function $t\mapsto t\ln t$ on $(0,1)$. This bound was obtained in the proof of Theorem 1.1 in [AGB21]. For the sake of completeness, we provide the proof as a separate lemma in Appendix B.
We claim that the integral on the left-hand side of the Brascamp–Lieb inequality is
@@
The inclusion $E\cap F\subset K\cap F\subset d\left(E\cap F\right)$ follows from Corollary 1.5, since $z$ from the assertion of the corollary is zero and $P_{F}E=E\cap F.$
@@ 5.5. Tightness of the results
The inclusions of Theorem 1.7 are optimal since they are already optimal for the John ellipsoid. The volume bounds of Lemma 5.7 and Lemma 5.8 coincide with the corresponding bounds for an $s$-dimensional section of a $d$-dimensional convex body in John position (see Theorem 1.1 of [AGB21]). In the centrally symmetric case $K=-K,$ we conjecture the following:
@@
Let an origin-centered ellipsoid $E$ be a solution to Problem 1 for a centrally symmetric convex body $K$ and a linear subspace $F$ in $\mathbb{R}^{d}$ of dimension $s\in\{0,1,\dots,d\}$. Then
@@ 6. Uniqueness of the solution
Consider the family of pairs $(S,x)$, where $F$ is a positive semidefinite $F$-operator, $x\in\mathbb{R}^{d}$ satisfying $x+S(K)\subseteq L$. We regard $(S,x)$ as a point of $\mathbb{R}^{d^{2}+d}$. It is easy to see that the family of pairs $(S,x)$ with the above condition is a compact, convex subset of $\mathbb{R}^{d^{2}+d}$, and thus, $\operatorname{vol}\nolimits_{d}\left(x+S(K)\right)=\det(S)\operatorname{vol}\nolimits_{d}K$ attains its maximum on it. Note also that if $x+S(K)$ maximizes $\operatorname{vol}\nolimits_{d}\left(x+S(K)\right)$ on this set, then $S$ is positive definite, as otherwise $\operatorname{vol}\nolimits_{d}\left(x+S(K)\right)=0$. Thus, there are some positive definite $F$-operators $S_{1},S_{2}$ and vectors $x_{1},x_{2}\in\mathbb{R}^{d}$ such that for $i=1,2$, $K_{i}=x_{i}+S_{i}(K)$.
By convexity, we have that $\bar{K}=\frac{x_{1}+x_{2}}{2}+\frac{1}{2}(S_{1}+S_{2})(K)\subseteq L$. On the other hand, it is well-known that for any two positive definite matrices $A,B$,
@@
with equality if and only if $A=B$ (see e.g. [HJ12, Corollary 7.6.8]). Since $\frac{1}{2}(S_{1}+S_{2})$ is also a positive definite $F$-operator, it follows from the maximality of volume that $S_{1}=S_{2}$, implying that $K_{2}$ and $K_{1}$ are translates of each other. ∎
@@
We note that it is not true that if $K_{1},K_{2}\in\mathcal{S}_{F}(K)$ are of maximal volume satisfying $K_{1},K_{2}\subseteq L$, then the translation vector moving $K_{1}$ to $K_{2}$ is necessarily parallel to ${F}^{\bot}$, as shown by the following example. Set $d=2$ and $K=\mathbf{B}^{2}$. We define $L$ as an $\mathbf{0}$-symmetric rectangle circumscribed about $\mathbf{B}^{2}$ such that the sides of $L$ are parallel to the lines $y=\pm x$, and the only points of $\operatorname{bd}L$ on $\mathbf{B}^{2}$ are the points ±$\pm\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right)$. Then an elementary computation shows that the only $\mathbf{0}$-symmetric $F$-ellipse contained in $L$ is $\mathbf{B}^{2}$, and this is of maximal area. On the other hand, moving this ellipse along the line $y=-x$ yields other maximal area $F$-ellipses in $L$.
@@ 7. $F$-revolution Löwner ellipsoids
    where $\Delta_{L}^{s}$ stands for the $s$-dimensional simplex inscribed in $\mathbf{B}^{s}.$
@@
    $P_{F}K\subset P_{F}E\subset d\cdot P_{F}K.$
@@
    the circumradius of $P_{{F}^{\bot}}K$ is at least $\sqrt{\frac{d-s}{d}},$
@@
    where $\Diamond^{s}$ stands for the standard cross-polytope.
@@
    $P_{F}K\subset P_{F}E\subset\sqrt{d}\cdot P_{F}K.$
@@ 7.2. Circumradius in ${F}^{\bot}$
Let $v_{1},\dots,v_{m}$ be vectors in $\mathbb{R}^{n}.$ Assume there are positive weights $\beta_{1},\dots,\beta_{m}$ such that
@@
7.3. Outer volume ratio in ${F}^{\bot}$
The proofs of our outer volume bounds follow from Theorem 1.2 of [AGB21], where the authors study lower bounds on the volume of projections of convex bodies in the Löwner position. Starting with the breakthrough work [Bar97], the main tool for obtaining outer volume ratio bounds is the following reverse form of the Brascamp–Lieb inequality:
@@ 7.3. Outer volume ratio in ${F}^{\bot}$
Let unit vectors $v_{1},\ldots,v_{m}\in\mathbb{R}^{n}$ and positive numbers $\beta_{1},\ldots,\beta_{m}$ satisfy $\sum\limits_{i\in[m]}\beta_{i}v_{i}\otimes v_{i}=\mathrm{Id}_{\mathbb{R}^{n}}.$ Define
$G(x)=\sup\left\{\prod\limits_{i\in[m]}g_{i}(\theta_{i})^{\beta_{i}}:\;x=\sum\limits_{i\in[m]}\beta_{i}\theta_{i}v_{i}\right\}$
for integrable functions $g_{1},\ldots,g_{m}\colon\mathbb{R}\to[0,\infty).$ Then,
$\int_{\mathbb{R}^{d}}G\geq\prod\limits_{i\in[m]}\left(\int g_{i}\right)^{\beta_{i}}.$
To simplify the reading, we provide a sketch of the proof, omitting technical details.
@@
Let $I\subset[m]$ be the set of those indices for which $P_{F}u_{i}\neq 0.$ Set $v_{i}=\frac{P_{F}u_{i}}{\left|P_{F}u_{i}\right|}$ and $\beta_{i}={\upalpha_{i}}{\left|P_{F}u_{i}\right|^{2}}$ for all $i\in I.$ Then
@@
We leave it to the reader to verify that the product of the integrals on the right-hand side of the reverse Brascamp–Lieb inequality (Lemma 7.6) equals $2^{s}\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}$. This is at least $2^{s}\left(\frac{d}{s}\right)^{\frac{s}{2}}$ by routine analysis.
@@
Following Barthe’s argument [Bar97], we lift the picture to $\widetilde{F}=F\times\mathbb{R}$, and write $(a,t)$ for a vector in $\widetilde{F}$ with $a\in F$ and $t\in\mathbb{R}.$ Let $I\subset[m]$ be the index set of those $i$ such that $P_{F}u_{i}\neq 0.$ Let $d^{\prime}=\sum\limits_{i\in[m]}\upalpha_{i}$, and define
@@
where we set $C(x)=+\infty$ if such a decomposition does not exist. Let $G(x)=e^{-C(x)}$.
@@
Applying Jensen’s inequality with $\sum\limits_{i\in I}\widetilde{\upalpha}_{i}\left|\widetilde{q}_{i}\right|^{2}=s+1$ and $\sum\limits_{i\in I}\widetilde{\upalpha}_{i}=d^{\prime}$, we obtain
@@
The bound is optimal whenever $s+1$ divides $d+1$. It is attained when $K=\Delta_{L}^{d}$ and $F$ is the subspace spanned by the centroids of pairwise disjoint $\left(\frac{d+1}{s+1}-1\right)$-dimensional faces of $\Delta_{L}^{d}.$
@@ Appendix A Tightness of the inradius bound
The goal of this section is to state and prove Theorem A.5 according to which the bound Theorem 1.7 (1) on the inradius is the best possible.
Define
@@
We choose vectors $v_{1},\dots,v_{n+1}$ that form a directionally dilated regular simplex along the direction of $a$.
For each $i\in[n]$, define the vector $p_{i}\in\mathbb{R}^{n}$ by
@@
Then $\left|p_{i}\right|^{2}=\frac{n-1}{n}.$ Define $q_{i}=\frac{p_{i}}{\left|p_{i}\right|}\in\mathbb{R}^{n}.$ Note that for each $i\in[n]$, the vectors $p_{i}$ and $q_{i}$ belong to ${a}^{\bot}$.
Define the vectors
@@
Now, fix a constant $\gamma$ satisfying
@@
$\gamma\sum_{i\in[n]}v_{i}+(1-n\gamma)v_{n+1}=\mathbf{0}.$
@@
Denote $\Delta=\operatorname{conv}\left\{v_{1},\dots,v_{n+1}\right\}$. Let us estimate the inradius of ${\Delta}^{\circ}.$
@@
Proof.
The vertex of ${\Delta}^{\circ}$ corresponding to the facet $\operatorname{conv}\{v_{1},\dots,v_{n}\}$ of $\Delta$ is
$x_{n+1}=\frac{1}{\sqrt{1-t^{2}}}a.$
Clearly, $\left\langle x_{n+1},v_{n+1}\right\rangle<0$ and $\left\langle x_{n+1},v_{i}\right\rangle=1$ for all $i\in[n]$.
By direct computation, for each $i\in[n]$,
$\left\langle x_{i},v_{i}\right\rangle=-\frac{1}{n\gamma}\cdot\frac{n-1}{t}\cdot t\left\langle q_{i},q_{i}\right\rangle-\frac{1-n\gamma}{n\gamma}\cdot\frac{1}{\sqrt{1-t^{2}}}\cdot\sqrt{1-t^{2}}\left\langle a,a\right\rangle=-\frac{n-n\gamma}{n\gamma},$
where we used the fact that $\left\langle q_{i},a\right\rangle=0$. Note that if inequality (A.1) holds, then $\left\langle x_{i},v_{i}\right\rangle<0$ for every $i\in[n]$.
Since $\left\langle q_{i},q_{j}\right\rangle=-\frac{1}{n-1}$ whenever $j\in[n]$ and $i\neq j$, we conclude that
$\left\langle x_{i},v_{j}\right\rangle=-\frac{1}{n\gamma}\cdot\frac{n-1}{t}\cdot t\left\langle q_{i},q_{j}\right\rangle-\frac{1-n\gamma}{n\gamma}\cdot\frac{1}{\sqrt{1-t^{2}}}\cdot\sqrt{1-t^{2}}\left\langle a,a\right\rangle=\frac{1}{n\gamma}-\frac{1-n\gamma}{n\gamma}=1$
for all $j\neq i$, as required for duality. Clearly, $\left\langle x_{i},v_{n+1}\right\rangle=1$ for all $i\in[n]$. The lemma follows. ∎
Remark A.2.
There is no mystery in the explicit formulas for the vertices of ${\Delta}^{\circ}$. By symmetry, the facet of ${\Delta}^{\circ}$ opposite to $x_{n+1}$ must be parallel to ${a}^{\bot}$. Therefore, the $a$-component can be found directly from the equation $\left\langle x_{i},v_{n+1}\right\rangle=1$. On the other hand, this facet must be a translation of the simplex $-\tau\operatorname{conv}\{q_{1},\dots,q_{n}\}$ for some positive $\tau$. Indeed, this facet is a homothet of ${\Delta}^{\circ}\cap{a}^{\bot}$ with center at $x_{n+1}$. The set ${\Delta}^{\circ}\cap{a}^{\bot}$ is the regular simplex
$\operatorname{conv}\left\{\frac{1-n}{t}q_{1},\dots,\frac{1-n}{t}q_{n}\right\},$
since ${\Delta}^{\circ}\cap{a}^{\bot}$ is the polar in ${a}^{\bot}$ of the projection of $\Delta$ onto ${a}^{\bot}$. This projection is the regular simplex $\operatorname{conv}\left\{tq_{1},\dots,tq_{n}\right\}$.
@@
Thus, $P_{L}\Delta$ is the triangle with vertices $v_{1}$, $v_{n+1}$, and $P_{L}v_{2}$. Hence, ${\Delta}^{\circ}\cap L$ is a triangle, say $T$. Let us compute its vertices. Since $x_{1}$ and $x_{n+1}$ lie in $L$, they are two of the vertices of $T$. The remaining vertex, say $y$, can be found using duality in $L$:
@@
In our setting, this shows that the incenter of ${\Delta}^{\circ}$ lies in $T$ and is the intersection point of the altitude $x_{n+1}h$ and the angle bisector from vertex $y$.
@@
From the right-angle triangle $x_{n+1}hy$,
@@
Since $yz$ is the angle bisector of triangle $yhx_{n+1}$, we have:
@@
Fix a positive integer $d$ such that $d>n\geq 2,$ and set $\gamma=\frac{1}{d}.$ Then, inequality (A.1) holds and the following hold as $t$ tends to $1^{-}$:
@@
In particular, for any positive integer $m>d-s$ and positive $\varepsilon,$ there is a configuration of $m$ vectors $b_{1},\dots,b_{m}$ of $\mathbb{R}^{d-s}$ and positive numbers $\beta_{1},\dots,\beta_{m}$ satisfying
Proof.
@@
Set $n=d-s.$ For each $i\in[n],$ set $b_{i}=\sqrt{\frac{\theta}{\sigma(t)}}v_{i}$ with $\theta=\frac{d-s}{d}.$ Set $b_{n+1}=\ldots=b_{m}=\sqrt{\frac{\theta}{\sigma(t)}}v_{n+1}.$ Then,
${\gamma\sum_{i\in[n]}b_{i}+\frac{(1-n\gamma)}{m-n}\sum_{i\in[m]\setminus[n]}b_{i}}=\sqrt{\frac{\theta}{\sigma(t)}}\left(\gamma\sum_{i\in[n]}v_{i}+(1-n\gamma)v_{n+1}\right)=\mathbf{0};$
@@
$\gamma\sum_{i\in[n]}\left|b_{i}\right|^{2}+\frac{(1-n\gamma)}{m-n}\sum_{i\in[m]\setminus[n]}\left|b_{i}\right|^{2}=\sigma(t).$
The inradius of the polar of $\operatorname{conv}\{b_{1}\dots,b_{m}\}$ is greater than that of the simplex $v_{1}\dots,v_{d-s+1}$ for $t$ sufficiently close to $1$ On the other hand, there is $t$ such that the latter is at least $(1-\varepsilon)\frac{d}{d-s}.$

### v2

@@ 5.2. Inradius in ${F}^{\bot}$
In the centrally symmetric case, we know that the origin is the center of an inball. Hence, the inradius cannot exceed the distance from the origin to a hyperplane $\left\{x\in{F}^{\bot}:\;\left\langle P_{{F}^{\bot}}u_{i},x\right\rangle=1\right\}$ in ${F}^{\bot}$ for some non-zero $P_{{F}^{\bot}}u_{i}.$ This distance is exactly $\frac{1}{\left|P_{{F}^{\bot}}u_{i}\right|}.$ Identities (1.1) and (1.3) imply that $\left|P_{{F}^{\bot}}u_{j}\right|^{2}\geq\frac{d-s}{d}$ for some $j\in[m].$ The bound follows. ∎
@@ 5.3. Volume ratio in $F$
(5.2) $\int_{\mathbb{R}^{n}}\prod_{i\in[m]}f_{i}(\left\langle x,v_{i}\right\rangle)^{\beta_{i}}\,\mathrm{d}x\leq\prod_{i\in[m]}\left(\int_{\mathbb{R}}f_{i}\right)^{\beta_{i}},$
@@
Let $I\subset[m]$ be the set of those indices $i\in[m]$ for which $P_{F}u_{i}\neq 0$. Set $v_{i}=\frac{P_{F}u_{i}}{\left|P_{F}u_{i}\right|}$ and $\beta_{i}={\upalpha_{i}}{\left|P_{F}u_{i}\right|^{2}}$ for all $i\in I.$ Then
@@
We leave it to the reader to check that the integral on the left-hand side of (5.2) is equal to the volume of $K_{\mathrm{J_{sym}}}\cap F$, as we identify $F$ with $\mathbb{R}^{n}$ (where $n=s$), and that the product of the integrals on the right-hand side of (5.2) is equal to $2^{s}\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{-\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}.$ The latter quantity is at most $2^{s}\left(\frac{d}{s}\right)^{\frac{s}{2}}$ by Jensen’s inequality, completing the proof of Lemma 5.7. ∎
@@
[…]
Let $I\subset[m]$ denote the subset of indices $i\in[m]$ for which $P_{F}u_{i}\neq 0$. Denote $d^{\prime}=\sum\limits_{i\in I}\upalpha_{i};$ and set
@@
We will denote the function $t\mapsto e^{-\frac{t}{\left|\widetilde{q}_{i}\right|}}\mathbf{1}_{[0,+\infty)}$ by $f_{i}$ for each $i\in I,$ where $\mathbf{1}_{S}$ stands for the indicator function of the set $S.$ Then, the product of the integrals on the right-hand side of (5.2) is equal to $\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}.$ The bound
(5.3) $\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}\leq\frac{(s+1)^{\frac{s+1}{2(d^{\prime}+1)}}}{(d^{\prime}+1)^{\frac{d^{\prime}+1}{2(d^{\prime}+1)}}}$
follows from the standard majorization argument for the convex function $t\mapsto t\ln t$ on (0,1)(0,1)( 0 , 1 ). This bound was obtained in the proof of Theorem 1.1 in [AGB21]. For the sake of completeness, we provide the proof as a separate lemma in Appendix B.
We claim that the integral on the left-hand side of (5.2) is
@@
The inclusion $E\cap F\subset K\cap F\subset d\left(E\cap F\right)$ follows from Theorem 1.5 with $z=\mathbf{0}$ and $P_{F}E=E\cap F\subset K\cap F.$
@@ 5.5. Tightness of the results
The inclusions of Theorem 1.7 are optimal since they are already optimal for the John ellipsoid as shown by the simplex and the cube. The volume bounds of Lemma 5.7 and Lemma 5.8 coincide with the corresponding bounds for an $s$-dimensional section of a $d$-dimensional convex body in John position (see Theorem 1.1 of [AGB21]). In the centrally symmetric case $K=-K,$ we conjecture the following:
@@
Let an origin-centered ellipsoid $E$ be a solution to Problem 1 for a centrally symmetric convex body $K$ and a linear subspace $F$ in $\mathbb{R}^{d}$ of dimension $s\in[d]$. Then
@@ 6. Uniqueness of the solution
[…]
Consider some positive definite $F$-operators $S_{1},S_{2}$ and vectors $x_{1},x_{2}\in\mathbb{R}^{d}$ such that for $i=1,2$, $K_{i}=x_{i}+S_{i}(K)$ maximize volume. By convexity, we have that $\bar{K}=\frac{x_{1}+x_{2}}{2}+\frac{1}{2}(S_{1}+S_{2})(K)\subseteq L$. On the other hand, it is well-known that for any two positive definite matrices $A,B$,
@@
with equality if and only if $A=B$ (see e.g. [HJ12, Corollary 7.6.8]). Since $\frac{1}{2}(S_{1}+S_{2})$ is also a positive definite $F$-operator, it follows from the maximality of volume that $S_{1}=S_{2}$, implying that $K_{1}$ and $K_{2}$ are translates of each other. ∎
@@
It is not true that if $K_{1},K_{2}\in\mathcal{S}_{F}(K)$ are of maximal volume satisfying $K_{1},K_{2}\subseteq L$, then the translation vector moving $K_{1}$ to $K_{2}$ is necessarily parallel to ${F}^{\bot}$, as shown by the following example. Set $d=2$ and $K=\mathbf{B}^{2}$. We define $L$ as an $\mathbf{0}$-symmetric rectangle containing $\mathbf{B}^{2}$ such that the sides of $L$ are parallel to the lines $y=\pm x$, and the only points of $\operatorname{bd}L$ on $\mathbf{B}^{2}$ are the points ±$\pm\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right)$. Let $F$ be the $y$-axis. Then an elementary computation shows that the only $\mathbf{0}$-symmetric $F$-ellipse inscribed in $L$ is $\mathbf{B}^{2}$, and thus, this is of maximal area. On the other hand, moving this ellipse along the line $y=-x$ yields other maximal area $F$-ellipses in $L$.
@@ 7. $F$-revolution Löwner ellipsoids
    where $\Delta_{L}^{s}$ stands for the regular $s$-dimensional simplex inscribed in $\mathbf{B}^{s}$;
@@
    $P_{F}K\subset P_{F}\mathbf{B}^{d}\subset d\cdot P_{F}K;$
@@
    the circumradius of $P_{{F}^{\bot}}K$ is at least $\sqrt{\frac{d-s}{d}}.$
@@
    where $\Diamond^{s}$ stands for the standard cross-polytope;
@@
    $P_{F}K\subset P_{F}\mathbf{B}^{d}\subset\sqrt{d}\cdot P_{F}K.$
@@ 7.2. Circumradius in ${F}^{\bot}$
Let $v_{1},\dots,v_{m}$ be vectors in $\mathbb{R}^{n}.$ Assume that there are positive weights $\beta_{1},\dots,\beta_{m}$ such that
@@
7.3. Outer volume ratio in $F$
The proofs of our outer volume bounds follow from Theorem 1.2 of [AGB21], where the authors study lower bounds on the volume of projections of convex bodies in the Löwner position. Starting with the breakthrough work [Bar97] of Barthe, the main tool for obtaining outer volume ratio bounds is the following reverse form of the Brascamp–Lieb inequality:
@@ 7.3. Outer volume ratio in $F$
Let unit vectors $v_{1},\ldots,v_{m}\in\mathbb{R}^{n}$ and positive numbers $\beta_{1},\ldots,\beta_{m}$ satisfy $\sum\limits_{i\in[m]}\beta_{i}v_{i}\otimes v_{i}=\mathrm{Id}_{\mathbb{R}^{n}}$, and $g_{1},\ldots,g_{m}\colon\mathbb{R}\to[0,\infty)$ be integrable functions. Define
$G(x)=\sup\left\{\prod\limits_{i\in[m]}g_{i}(\theta_{i})^{\beta_{i}}:\;\theta_{i}\in\mathbb{R}\text{ and }x=\sum\limits_{i\in[m]}\beta_{i}\theta_{i}v_{i}\right\}.$
Then,
(7.1) $\int_{\mathbb{R}^{n}}G\geq\prod\limits_{i\in[m]}\left(\int g_{i}\right)^{\beta_{i}}.$
To simplify the reading, we provide a sketch of the proof of the following lemma, omitting technical details.
@@
Let $I\subset[m]$ be the set of those indices $i\in[m]$ for which $P_{F}u_{i}\neq 0.$ Set $v_{i}=\frac{P_{F}u_{i}}{\left|P_{F}u_{i}\right|}$ and $\beta_{i}={\upalpha_{i}}{\left|P_{F}u_{i}\right|^{2}}$ for all $i\in I.$ Then
@@
We leave it to the reader to verify that the product of the integrals on the right-hand side of (7.1) equals $2^{s}\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}$. This is at least $2^{s}\left(\frac{d}{s}\right)^{\frac{s}{2}}$ by Jensen’s inequality.
@@
[…]
@@
where we set $C(x)=+\infty$, if such a decomposition does not exist. Let $G(x)=e^{-C(x)}$.
@@
Applying Jensen’s inequality with $\sum\limits_{i\in I}\widetilde{\upalpha}_{i}\left|\widetilde{q}_{i}\right|^{2}=s+1$ and $\sum\limits_{i\in I}\widetilde{\upalpha}_{i}=d^{\prime}+1$, we obtain
@@
The bound is optimal whenever $s+1$ divides $d+1$. It is attained when $K=\Delta_{L}^{d}$ and $F$ is the subspace spanned by the centroids of $s+1$ pairwise disjoint $\left(\frac{d+1}{s+1}-1\right)$-dimensional faces of $\Delta_{L}^{d}.$
@@ Appendix A Tightness of the inradius bound
The goal of this section is to state and prove Theorem A.4 according to which the bound given in (1) of Theorem 1.7 on the inradius is the best possible.
Set
@@
We will define vectors $v_{1},\dots,v_{n+1}$ that form a directionally dilated regular simplex along the direction of $a$.
For each $i\in[n]$, define the vector $p_{i}\in\mathbb{R}^{n}$ whose $j$-th coordinate is given by
@@
Then $\left|p_{i}\right|^{2}=\frac{n-1}{n}.$ Define $q_{i}=\frac{p_{i}}{\left|p_{i}\right|}\in\mathbb{R}^{n}.$ Note that for each $i\in[n]$, the vectors $p_{i}$ and $q_{i}$ belong to ${a}^{\bot}$, and that $\left\langle q_{i},q_{j}\right\rangle=-\frac{1}{n-1}$ for any $1\leq i<j\leq n$.
For any $t\in(0,1)$, define the vectors
@@
Now, fix a positive constant $\gamma$ satisfying
@@
(A.2) $\gamma\sum_{i\in[n]}v_{i}+(1-n\gamma)v_{n+1}=\mathbf{0}.$
@@
Denote $\Delta=\operatorname{conv}\left\{v_{1},\dots,v_{n+1}\right\}$. Let us estimate the inradius of the polar ${\Delta}^{\circ}$ of $\Delta$. A direct computation yields the following:
@@
Thus, $P_{L}\Delta$ is the triangle with vertices $v_{1}$, $v_{n+1}$, and $P_{L}v_{2}$. Hence, ${\Delta}^{\circ}\cap L$ is a triangle, say $T$. Let us compute its vertices. Since $x_{1}$ and $x_{n+1}$ lie in $L$, they are two of the vertices of $T$. The remaining vertex $y$ can be found using duality in $L$:
@@
In our setting, this shows that the incenter of ${\Delta}^{\circ}$ lies in $T$ and is the intersection point $z$ of the altitude $x_{n+1}h$ and the angle bisector from vertex $y$.
@@
From the right triangle $x_{n+1}hy$,
@@
Since $yz$ is an angle bisector of the triangle $yhx_{n+1}$, we have:
@@
Fix non-negative integers $d,s$ and $m$ with $m>d-s\geq 2$ and a positive $\varepsilon$. Then there is a configuration of $m$ vectors $b_{1},\dots,b_{m}$ in $\mathbb{R}^{d-s}$ and positive numbers $\beta_{1},\dots,\beta_{m}$ satisfying
Proof.
Set $\gamma=\frac{1}{d}$ and $n=d-s$. Then, with the notation introduced previously, inequality (A.1) holds and the following hold as $t$ tends to $1^{-}$:
@@
For each $i\in[n],$ set $b_{i}=\sqrt{\frac{\theta}{\sigma(t)}}v_{i}$ with $\theta=\frac{d-s}{d}.$ Set $b_{n+1}=\ldots=b_{m}=\sqrt{\frac{\theta}{\sigma(t)}}v_{n+1}$. Also set $\beta_{i}=\gamma$ for $i\in[n]$ and $\beta_{i}=\frac{1-n\gamma}{m-n}$ for $i\in[m]\setminus[n]$. Then clearly, $\sum\limits_{i\in[m]}\beta_{i}=1$. Moreover,
[…]
@@
$\sum\limits_{i\in[m]}\beta_{i}\left|b_{i}\right|^{2}=\gamma\sum_{i\in[n]}\left|b_{i}\right|^{2}+\frac{(1-n\gamma)}{m-n}\sum_{i\in[m]\setminus[n]}\left|b_{i}\right|^{2}=\theta.$
The inradius of the polar of $\operatorname{conv}\{b_{1},\dots,b_{m}\}$ is greater than that of the simplex $v_{1}\dots v_{d-s+1}$ for $t$ sufficiently close to 1.1.1 . On the other hand, there is a $t$ such that the latter is at least $(1-\varepsilon)\frac{d}{d-s}.$